A total of 14000 dollars is invested in two simple interest baring accounts. One acccount has an interest rate of 8% and the other a rte of 5.5%. The amount of interest earned for 1 year was 982.50. How much was invested in each account?
Answer by checkley77(12569)

.08x+.055(14,000-x)=982.50
.08x+770-.055x=982.50
.025x=982.50-770
.025x=212.5
x=212.50/.025
x=8,500 amount invested @ 8%.
14,000-8,500=5,500 amount invested @ 5.5%
Proof:
.08*8,500+.055*5,500=982.50
680+302.50=982.50
982.50=982.50